You are 12.5 meters away from a 2-story residential building with an approximate height of 7 meters. At the same time, you are 21 meters away from the church, how high is the church? Provide an illustration and a Solution
Found 3 solutions by Theo, math_tutor2020, ikleyn:
Answer by Theo(13342)

if you assume that the church is in the same line as the 2 story residential building and that the angle of elevation from you to the building is the same angle of elevation from you to the church, then you have 2 similar right triangles that have equal angles and have proportional sides.

here's my diagram.

ddddd

the angle of elevation is T.

tan(T) = 7/12.5 and tan(T) = x/21.

since they are both equal to tan(T), then they are equal to each other and you get 7/12.5 = x/21.
cross multiply to get 7*21 = 12.5*x
solve for x to get x = 7*21/12.5 = 11.76.

11.76 is the height of the church.

T is equal to arctan(7/12.5) and arctan(11.76/21).

use your calculator to find that T = 29.24882634 degrees in both cases.

here's my diagram.





Answer by math_tutor2020(3817)


The wording of your question is a bit vague.
If tutor @Theo has the correct diagram, then follow the steps shown below:
x/7 = 21/12.5
12.5x = 7*21
x = 7*21/12.5
x = 11.76

The church would have an appoximate height of roughly 11.76 meters.

The initial ratio in the 1st step is valid because of the similar triangles.

Answer by ikleyn(52781)

.

For the future generations of students, who will read this post.

The problem's formulation in the post is  FATALLY  incomplete,
since it does not have a key description,  connecting the two objects.

So,  do not consider it seriously.

And never formulate your problems this way.


Respect yourself and respect the tutors.